OpenStudy (lilshane):

State whether every point on the graph is a solution of the equation. Explain why or why not. 3x – 2y = –6

OpenStudy (lilshane):

A. Every point shown on the graph is a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is true for all the points. B. Every point shown on the graph is not a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is not true for all the points.

OpenStudy (lilshane):

@iGreen @answermyquestions

OpenStudy (texaschic101):

kinda need the graph to see the points

OpenStudy (texaschic101):

ok...our points are : (2,5) , (0,3) , (-2,0) , (-2,-4) now we will sub into the equation to see if they are true 3x - 2y = 6 using points (2,5)...x = 2 and y = 5 3(2) - 2(5) = 6 6 - 10 = 6 -4 = 6 ---no way this is true next one.. 3x - 2y = 6 using points (0,3)...x = 0 and y = 3 3(0) - 2(3) = 6 0 - 6 = 6 -6 = 6 -- not true either and if you subbed in your next set of points, you will find that they are incorrect. understand ?

OpenStudy (lilshane):

So B?

OpenStudy (texaschic101):

yep...its B

OpenStudy (lilshane):

Thanks!

OpenStudy (lilshane):

Can you help with some more?

OpenStudy (texaschic101):

no problem...happy to help :)

OpenStudy (texaschic101):

I can try

OpenStudy (lilshane):

How many solutions are there to the equation y = 14x – 9? Explain your answer. A. There are no solutions because no value can be used for the variable x for the equation to remain true. B. There is one solution because only one value can be used for the variable x for the equation to remain true. C. There are 2 solutions because only 2 values can be used for the variables for the equation to remain true. D. There are an infinite number of solutions because any value can be used for the variable x and the equation would remain true.

OpenStudy (texaschic101):

lets check.....lets sub in 1 for x y = 14x - 9 y = 14(1) - 9 y = 14 - 9 y = 5 so when x is 1, then y = 5 now sub in 2 for x y = 14x - 9 y = 14(2) - 9 y = 28 - 9 y = 19 so when x is 2, then y = 19 I would go with D

OpenStudy (texaschic101):

no matter what number you sub in for x, there is always gonna be a y